# literature.bib

### v2

@@ 6 Problem Variants with Continuous Variables
If the item sets are not disjoint, the situation is different. In fact, we might obtain a bilevel optimization problem whose optimal value is not attained. This is an interesting property that (mixed-integer) bilevel optimization problems can have, even if the corresponding single-level problems do not indicate such a behavior. The reason here is that the binary follower’s problem might introduce an unnatural discontinuity into the leader’s objective. We refer to, e.g., [vicentesavardjudice1996] for more examples of such a behavior in bilevel optimization problems with continuous leader’s variables and integer follower’s variables.
@@ 6.2 The Robust Continuous Bilevel Selection Problem
We now extend the general structural observations of Section 6.1 to the corresponding robust problem. We will first describe some general insights and show in Example 26 that the Robust Continuous Bilevel Selection Problem is not equivalent to the binary version, in contrast to the problem without uncertainty. Afterwards, we will investigate the Robust Continuous Bilevel Selection Problem for specific types of uncertainty sets: In Sections 6.2.1, 6.2.2, and 6.2.3, we give polynomial-time algorithms for the Robust Continuous Bilevel Selection Problem with discrete uncertainty, interval uncertainty, and discrete uncorrelated uncertainty, respectively. While the results for discrete and interval uncertainty make use of similar ideas as before and in [buchheimhenke2022], a modified approach is necessary in case of discrete uncorrelated uncertainty. The latter also provides a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem with discrete uncorrelated uncertainty, which answers an open question stated in [buchheimhenke2022]; see Section 6.2.3.1.
@@ 6.2.3 Discrete Uncorrelated Uncertainty
Nevertheless, one can solve the Robust Continuous Bilevel Selection Problem and its adversary’s problem in case of discrete uncorrelated uncertainty in polynomial time using similar ideas as for interval uncertainty, as we show in this section. Moreover, in Section 6.2.3.1 we present a generalization of the ideas, resulting in pseudopolynomial-time algorithms for the adversary’s problem and the leader’s problem of the Robust Bilevel Continuous Knapsack Problem with discrete uncorrelated uncertainty, adding to the complexity results in [buchheimhenke2022].
@@ 18                  $Y_{e^{*},\delta^{*}}:=\mathcal{E}_{e^{*},\delta^{*}}^{-}\cup Y_{e^{*},\delta^{*}}^{0}$
Remark 32.
In Sections 4.2 and 6.2.2, for interval uncertainty, we assumed that there are no one-point intersections between the intervals in the uncertainty set; see Remark 2 in [buchheimhenke2022] for more details. A corresponding assumption here, which ensures that distinguishing the optimistic and the pessimistic case is not necessary, is that the sets $\mathcal{U}_{e}$ are pairwise disjoint. However, analogously to interval uncertainty, Algorithm 5 is also correct for the pessimistic setting without this assumption, while an adjusted definition of the item sets $\mathcal{E}_{e^{*},\delta^{*}}^{-}$ and $\mathcal{E}_{e^{*},\delta^{*}}^{0}$ is required for the optimistic setting, similarly to Remark 2 in [buchheimhenke2022].
@@
7for $e\in{\mathcal{E}_{f}}$ do
@@ 18                        $Y_{e^{*},\delta^{*}}:=\mathcal{E}_{e^{*},\delta^{*}}^{-}\cup Y_{e^{*},\delta^{*}}^{0}$
24
25join all $g_{f}^{{b_{f}^{*}}}$ (with range $[b-{b_{f}^{*}}-1,b-{b_{f}^{*}}]$) to a function ${g_{f}}$ with range $[{b_{l}^{-}},{b_{l}^{+}}]$
26compute $g:={g_{l}}+{g_{f}}$
27compute some ${b_{l}^{*}}\in\operatorname*{argmin}\{g({b_{l}})\mid{b_{l}}\in[{b_{l}^{-}},{b_{l}^{+}}]\}$
@@
Algorithm 6 solves the Robust Continuous Bilevel Selection Problem with a discrete uncorrelated uncertainty set $\mathcal{U}=\prod_{e\in{\mathcal{E}_{f}}}\mathcal{U}_{e}$ in time $O(nu\log(u))$, where $u=\sum_{e\in{\mathcal{E}_{f}}}\lvert\mathcal{U}_{e}\rvert$.
We emphasize that the term $u$ and therefore the running times stated in Theorems 31 and 33 are polynomial in the input size, while the total size $\lvert\mathcal{U}\rvert$ of the discrete uncorrelated uncertainty set is exponential in the input size. Moreover, if the uncertainty set is given by sets $\mathcal{U}_{e}$ of constant size, in particular if each $\mathcal{U}_{e}$ consists of only two values $d^{-}(e)$ and $d^{+}(e)$, then the running time of Algorithm 5 is $O(n^{2})$, and the running time of Algorithm 6 is $O(n^{2}\log n)$, like in case of interval uncertainty; see Theorems 8 and 29, Corollary 14 and Theorem 30.
@@
In the corresponding versions of the Robust Bilevel Continuous Knapsack Problem, the difference between the two approaches is more significant: The Continuous Knapsack Problem can be solved in polynomial time, while the binary Knapsack Problem cannot. Therefore, the Robust Bilevel Continuous Knapsack Problem with interval uncertainty can be solved in polynomial time only by applying the strategy that solves a Continuous Knapsack Problem in each iteration; see [buchheimhenke2022]. In case of discrete uncorrelated uncertainty, this polynomial-time approach does not work, analogously to the discussion in the beginning of Section 6.2.3. In fact, the Robust Bilevel Continuous Knapsack Problem has been shown to be weakly NP-hard in [buchheimhenke2022]. However, the new approach we developed in this section can be generalized to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem; see Section 6.2.3.1.
6.2.3.1 A Pseudopolynomial-Time Algorithm for the Robust Bilevel Continuous Knapsack Problem
We now adapt the algorithms for the Robust Continuous Bilevel Selection Problem with discrete uncorrelated uncertainty that we presented above, in order to apply them to the related Robust Bilevel Continuous Knapsack Problem from [buchheimhenke2022]. As already indicated in Remark 34, this will lead to the first pseudopolynomial-time algorithm for this problem, while its weak NP-hardness has been proved in Theorem 3 of [buchheimhenke2022]. Also the adversary’s problem is weakly NP-hard (see Theorem 4 of [buchheimhenke2022]), and the generalization of Algorithm 5 gives a pseudopolynomial-time algorithm for it as well.
@@ 6.2.3.1 A Pseudopolynomial-Time Algorithm for the Robust Bilevel Continuous Knapsack Problem
Secondly, when fixing the fractional item of the follower’s Continuous Knapsack Problem, the fraction with which it is selected is not uniquely determined by the capacity $b$, in contrast to the case where the follower solves a Continuous Selection Problem. If $b$ is not an integer, in the latter case, the fraction is always given by $b-\lfloor b\rfloor$, and we know that $b$ items are selected before the fractional item. We made use of this property in Algorithm 5. In the more general case of a Continuous Knapsack Problem, the items selected before the fractional item $e^{*}$ are known to have an integer total size as well (as we assume all item sizes to be integers), but this can be any integer number $b^{*}$ from $\lfloor b-a(e^{*})\rfloor+1$ to $\lfloor b\rfloor$ now, corresponding to fractions $y_{e^{*}}=\frac{1}{a(e)}(i+(b-\lfloor b\rfloor))$ for $i\in\{0,\dots,a(e)-1\}$ (in reverse order). Therefore, we enumerate all these possibilities for $b^{*}$ in addition to the possible fractional items $e^{*}$ and their follower’s item values $\delta^{*}$ in Algorithm 7. Note that Line 7 of Algorithm 7 combines this with the capacity condition that is used also in previous versions of the algorithm; see, e.g., Line 5 of Algorithm 5. When solving the leader’s problem, we have to enumerate all reasonable integer capacities $b^{*}$ anyway (see Algorithm 6) and do this in a similar way also in Algorithm 8.
@@ 10             $\mathcal{E}_{e^{*},\delta^{*}}^{-}:=\{e\in\mathcal{E}\mid d^{-}(e)/a(e)>\delta^{*}/a(e^{*})\}$
Using the same arguments as in [buchheimhenke2022] and Section 6.2.3 above, one can see that the function $g$ computed by Algorithm 8 is exactly the leader’s objective function for the Robust Bilevel Continuous Knapsack Problem. The algorithm’s running time can be estimated similarly to Theorems 33 and 35. Note that the term $a(\mathcal{E})$ enters the running time due to the subroutine solving a binary Knapsack Problem as well as due to the enumeration of $b^{*}$. Together, we get:
@@
7 Conclusion
In relation to [buchheimhenke2022], we emphasize the following insights regarding the complexity of the Robust Bilevel Continuous Knapsack Problem and the Robust Continuous Bilevel Selection Problem, which is simpler in terms of the items not having different sizes. In case of discrete uncorrelated uncertainty, the leader’s as well as the adversary’s problem have been shown to be NP-hard in [buchheimhenke2022], while the corresponding Robust Continuous Bilevel Selection Problem is still solvable in polynomial time. In Section 6.2.3.1, we generalized our new approach to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem and thus showed that the problem with discrete uncorrelated uncertainty is in fact only weakly NP-hard, but not strongly NP-hard, answering an open question stated in [buchheimhenke2022].
Another interesting observation is that the variant of the Bilevel Selection Problem with continuous leader’s variables and binary follower’s variables yields a bilevel optimization problem whose optimal value may not be attained (see Example 25). This demonstrates that, in general, bilevel optimization problems containing integer and continuous variables should be handled with care in this respect. It could be interesting to study this problem variant further in view of computing feasible solutions that approach the optimal value.

### v3

@@ 6 Problem Variants with Continuous Variables
If the item sets are not disjoint, the situation is different. In fact, we might obtain a bilevel optimization problem whose optimal value is not attained; see Example 25. This is an interesting property that (mixed-integer) bilevel optimization problems can have, even if the corresponding single-level problems do not indicate such a behavior. The reason here is that the binary follower’s problem might introduce an unnatural discontinuity into the leader’s objective. We refer to, e.g., [vicentesavardjudice1996] for more examples of such a behavior in bilevel optimization problems with continuous leader’s variables and integer follower’s variables.
@@ 6.2 The Robust Continuous Bilevel Selection Problem
We now extend the general structural observations of Section 6.1 to the corresponding robust problem. We will first describe some general insights and show in Example 26 that the Robust Continuous Bilevel Selection Problem is not equivalent to the binary version, in contrast to the problem without uncertainty. Afterwards, we will investigate the Robust Continuous Bilevel Selection Problem for specific types of uncertainty sets: In Sections 6.2.1, 6.2.2, and 6.2.3, we give polynomial-time algorithms for the Robust Continuous Bilevel Selection Problem with discrete uncertainty, interval uncertainty, and discrete uncorrelated uncertainty, respectively. While the results for discrete and interval uncertainty make use of similar ideas as before and in [buchheimhenke2022], a modified approach is necessary in case of discrete uncorrelated uncertainty. The latter also provides a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem with discrete uncorrelated uncertainty, which answers an open question stated in [buchheimhenke2022]; see Appendix A.
@@ 6.2.3 Discrete Uncorrelated Uncertainty
Nevertheless, one can solve the Robust Continuous Bilevel Selection Problem and its adversary’s problem in case of discrete uncorrelated uncertainty in polynomial time using similar ideas as for interval uncertainty, as we show in this section. Moreover, in Appendix A we present a generalization of the ideas, resulting in pseudopolynomial-time algorithms for the adversary’s problem and the leader’s problem of the Robust Bilevel Continuous Knapsack Problem with discrete uncorrelated uncertainty, adding to the complexity results in [buchheimhenke2022].
@@ 18                  $Y_{e^{*},\delta^{*}}:=\mathcal{E}_{e^{*},\delta^{*}}^{-}\cup Y_{e^{*},\delta^{*}}^{0}$
6if ${b_{l}^{-}}={b_{l}^{+}}$ then
7       ${b_{l}^{*}}:={b_{l}^{-}}$
8else
9       compute ${g_{l}}:=\textsc{PLF}({\mathcal{E}_{l}},c\restriction{\mathcal{E}_{l}},{b_{l}^{-}},{b_{l}^{+}})$
10      for $e\in{\mathcal{E}_{f}}$ do
@@ 21                              $Y_{e^{*},\delta^{*}}:=\mathcal{E}_{e^{*},\delta^{*}}^{-}\cup Y_{e^{*},\delta^{*}}^{0}$
27
28      join all $g_{f}^{{b_{f}^{*}}}$ (with range $[b-{b_{f}^{*}}-1,b-{b_{f}^{*}}]$) to a function ${g_{f}}$ with range $[{b_{l}^{-}},{b_{l}^{+}}]$
29      compute $g:={g_{l}}+{g_{f}}$
30      compute some ${b_{l}^{*}}\in\operatorname*{argmin}\{g({b_{l}})\mid{b_{l}}\in[{b_{l}^{-}},{b_{l}^{+}}]\}$
31
@@
Algorithm 6 solves the Robust Continuous Bilevel Selection Problem with a discrete uncorrelated uncertainty set $\mathcal{U}=\prod_{e\in{\mathcal{E}_{f}}}\mathcal{U}_{e}$ in time $O(nu\log(nu))$, where $u=\sum_{e\in{\mathcal{E}_{f}}}\lvert\mathcal{U}_{e}\rvert$.
We emphasize that the term $u$ and therefore the running times stated in Theorems 31 and 32 are polynomial in the input size, while the total size $\lvert\mathcal{U}\rvert$ of the discrete uncorrelated uncertainty set is exponential in the input size. Moreover, if the uncertainty set is given by sets $\mathcal{U}_{e}$ of constant size, in particular if each $\mathcal{U}_{e}$ consists of only two values $d^{-}(e)$ and $d^{+}(e)$, then the running time of Algorithm 5 is $O(n^{2})$, and the running time of Algorithm 6 is $O(n^{2}\log n)$, like in case of interval uncertainty; see Theorems 8 and 29, Corollary 14 and Theorem 30.
@@
In the corresponding versions of the Robust Bilevel Continuous Knapsack Problem, the difference between the two approaches is more significant: The Continuous Knapsack Problem can be solved in polynomial time, while the binary Knapsack Problem cannot. Therefore, the Robust Bilevel Continuous Knapsack Problem with interval uncertainty can be solved in polynomial time only by applying the strategy that solves a Continuous Knapsack Problem in each iteration; see [buchheimhenke2022]. In case of discrete uncorrelated uncertainty, this polynomial-time approach does not work, analogously to the discussion in the beginning of Section 6.2.3. In fact, the Robust Bilevel Continuous Knapsack Problem has been shown to be weakly NP-hard in [buchheimhenke2022]. However, the new approach we developed in this section can be generalized to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem; see Appendix A.
Remark 34.
In Sections 4.2 and 6.2.2, for interval uncertainty, we assumed that there are no one-point intersections between the intervals in the uncertainty set, in order to avoid discussing technical details related to the optimistic and the pessimistic case. For discrete uncorrelated uncertainty, one could assume that the sets $\mathcal{U}_{e}$ are pairwise disjoint. However, analogously to interval uncertainty, Algorithms 5 and 6, as they are stated above, are also correct for the pessimistic setting without this assumption, while an adjusted definition of the item sets $\mathcal{E}_{e^{*},\delta^{*}}^{-}$ and $\mathcal{E}_{e^{*},\delta^{*}}^{0}$ is required for the optimistic setting, similarly to Remark 2 in [buchheimhenke2022].
7 Conclusion
In relation to [buchheimhenke2022], we emphasize the following insights regarding the complexity of the Robust Bilevel Continuous Knapsack Problem and the Robust Continuous Bilevel Selection Problem, which is simpler in terms of the items not having different sizes. In case of discrete uncorrelated uncertainty, the leader’s as well as the adversary’s problem have been shown to be NP-hard in [buchheimhenke2022], while the corresponding Robust Continuous Bilevel Selection Problem is still solvable in polynomial time. In Appendix A, we generalize our new approach to a pseudopolynomial-time algorithm for the Robust Bilevel Continuous Knapsack Problem and thus showed that the problem with discrete uncorrelated uncertainty is in fact only weakly NP-hard, but not strongly NP-hard, answering an open question stated in [buchheimhenke2022].
Another interesting observation is that the variant of the Bilevel Selection Problem with continuous leader’s variables and binary follower’s variables yields a bilevel optimization problem whose optimal value may not be attained (see Example 25). This demonstrates that, in general, bilevel optimization problems containing integer and continuous variables should be handled with care in this respect. It could be interesting to study this problem variant further, e.g., in view of computing the optimal value or computing feasible solutions that approach the optimal value.
Appendix A Appendix: A Pseudopolynomial-Time Algorithm for the Robust Bilevel Continuous Knapsack Problem
We now adapt the algorithms for the Robust Continuous Bilevel Selection Problem with discrete uncorrelated uncertainty that we presented in Section 6.2.3, in order to apply them to the related Robust Bilevel Continuous Knapsack Problem from [buchheimhenke2022]. As indicated in Remark 33, this will lead to the first pseudopolynomial-time algorithm for this problem, while its weak NP-hardness has been proved in Theorem 3 of [buchheimhenke2022]. Also the adversary’s problem is weakly NP-hard (see Theorem 4 of [buchheimhenke2022]), and we give a pseudopolynomial-time algorithm for it as well.
@@ Appendix A Appendix: A Pseudopolynomial-Time Algorithm for the Robust Bilevel Continuous Knapsack Problem
Secondly, when fixing the fractional item of the follower’s Continuous Knapsack Problem, the fraction with which it is selected is not uniquely determined by the capacity $b$, in contrast to the case where the follower solves a Continuous Selection Problem. If $b$ is not an integer, in the latter case, the fraction is always given by $b-\lfloor b\rfloor$, and we know that $b$ items are selected before the fractional item. We made use of this property in Algorithm 5. In the more general case of a Continuous Knapsack Problem, the items selected before the fractional item $e^{*}$ are known to have an integer total size as well (as we assume all item sizes to be integers), but this can be any integer number $b^{*}$ from $\lfloor b-a(e^{*})\rfloor+1$ to $\lfloor b\rfloor$ now, corresponding to fractions $y_{e^{*}}=\frac{1}{a(e^{*})}(i+(b-\lfloor b\rfloor))$ for $i\in\{0,\dots,a(e^{*})-1\}$ (in reverse order). Therefore, we enumerate all these possibilities for $b^{*}$ in addition to the possible fractional items $e^{*}$ and their follower’s item values $\delta^{*}$ in Algorithm 7. Note that Line 7 of Algorithm 7 combines this with the capacity condition that is used also in previous versions of the algorithm; see, e.g., Line 5 of Algorithm 5. When solving the leader’s problem, we have to enumerate all reasonable integer capacities $b^{*}$ anyway (see Algorithm 6) and do this in a similar way also in Algorithm 8.
@@ 10             $\mathcal{E}_{e^{*},\delta^{*}}^{-}:=\{e\in\mathcal{E}\mid d^{-}(e)/a(e)>\delta^{*}/a(e^{*})\}$
Using the same arguments as in [buchheimhenke2022] and Section 6.2.3, one can see that the function $g$ computed by Algorithm 8 is exactly the leader’s objective function for the Robust Bilevel Continuous Knapsack Problem. The algorithm’s running time can be estimated similarly to Theorems 32 and 35. Note that the term $a(\mathcal{E})$ enters the running time due to the subroutine solving a binary Knapsack Problem as well as due to the enumeration of $b^{*}$. Together, we get: